eddies age is 4 times marco's.in 10 years eddie will be twice as ikd as marco will be.how old is each now.thanks alot for your time and help
Answer by JimboP1977(311)

E = Eddie's current age
M = Marco's current age
E/4 = M (1)
After 10 years Eddie is twice Marco's age:


(E+10)/2 = M + 10 (2)

Make (1) = (2) by adding 10 to equation (1) :
E/4 + 10 = M + 10 = (E+10)/2
so
E/4 + 10 = (E+10)/2
Rearrange to get E on its own:
10 = (E+10)/2 - E/4
10 = (2E+20)/4 - E/4
10 = (E+20)/4
40-20 = E
So Eddie's original age was 20 and Marco's was 5. So their new age after 10 years is 30 and 15 respectively.
Is this clear?
Does it help?